OpenStudy (antiprime):

I drive at an average speed of 30 miles per hour to the railroad station each morning and just catch my train. On a particular morning there was a lot of traffic and at the halfway point I found I had averaged only 15 miles per hour. How fast must I drive for the rest of the way to catch my train?

OpenStudy (antiprime):

i think i know the anwser

OpenStudy (welshfella):

what do you think it is?

OpenStudy (antiprime):

The train is just about to leave the station and there is no way I will be able to catch it this time. '

ganeshie8 (ganeshie8):

this is really a very interesting question !

OpenStudy (antiprime):

yes ik and its math related

OpenStudy (igreen):

You would have to go double the speed since you were only going half the speed from half way..right?

ganeshie8 (ganeshie8):

basically you're trying to find \(v\) that satisfies below equation : \[\dfrac{x/2}{15} + \dfrac{x/2}{v} = \dfrac{x}{30}\]

OpenStudy (antiprime):

yes but on original days he almost misses it

ganeshie8 (ganeshie8):

solving gives \(v = \infty\)

OpenStudy (antiprime):

and u cant achive that speed

OpenStudy (igreen):

If the car is like the car you have in your picture, you shouldn't have to worry about speed :P

OpenStudy (antiprime):

lol @iGreen

ganeshie8 (ganeshie8):

you can still catch the train if you know special relativity

OpenStudy (antiprime):

well we are assumig the person doesnt

ganeshie8 (ganeshie8):

you can go at some appropriate speed such that the remaining distance shrinks by the required amount so that you catch the train at exact time

OpenStudy (antiprime):

how can he avrage 30 mph tho

ganeshie8 (ganeshie8):

heard of space contraction and time dilation before ?

ganeshie8 (ganeshie8):

hey we cannot achieve if we stick to old newtonian physics i think we will have some hope if we use special relativity..

OpenStudy (antiprime):

kinda

OpenStudy (antiprime):

this is still complicated

ganeshie8 (ganeshie8):

the simple answer to you question is impossible yeah :)

OpenStudy (antiprime):

if u can time warp then mabye lol

ganeshie8 (ganeshie8):

space contraction doesn't help here as your velocity relative to earth in radial direction is 0

OpenStudy (antiprime):

true

ganeshie8 (ganeshie8):

oh no, it should still apply...

OpenStudy (antiprime):

@welshfella it still can happen some way

OpenStudy (anonymous):

this question doesn't need 12 people answering it

OpenStudy (anonymous):

try helping someone else

OpenStudy (welshfella):

yes? perhaps thers something wrong with my calculations then

OpenStudy (antiprime):

it still can happen if we have the right automobile

OpenStudy (welshfella):

ah yes the Batmobile!!!!

OpenStudy (antiprime):

anyways thxs guys for helping i gtg <3 U GUYS

OpenStudy (welshfella):

yw - interesting question...

ganeshie8 (ganeshie8):

you can't catch the train because it has already started and traveling to past is impossible as far as we know

OpenStudy (mathmale):

If you seriously want to solve this problem (or to prove that it can't be solved), please represent your unknown quantities. Regardless of the speed traveled, the distance d from home to station is a constant. The commonly used formula for distance is d=rt, where r is the rate of travel (and could be the average rate of travel). Is there any particular speed that in combination with 15 mph would result in this commuter's reaching the station just in time to catch the train?